# Understanding the Performance Gap in Preference Learning: A Dichotomy of RLHF and DPO

### v1

@@ C.3 Proof of Proposition 3
$\mathcal{L}_{\textup{DPO}}(\theta)$ $=-\sum_{a\neq a^{\prime}}\left[\sigma(r^{\star}(a)-r^{\star}(a^{\prime}))\log\sigma\left(\beta\,\theta^{\top}(\psi(a)-\psi(a^{\prime})\right)\right]$
@@ C.4 Proof of Proposition 5
For the strict gap, consider a multi-armed bandit setting analogous to Section C.3: first, define the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
@@
The key difference from the earlier construction lies in the choice of model classes. We define: the linear reward class =$\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d}\}$ by $r_{\phi}(a):=\phi^{\top}\psi(a)$, and the policy class $\Pi=\Delta(\mathcal{Y})$ with reference policy $\pi_{{\mathsf{ref}}}=\textrm{Unif}(\mathcal{Y})$. This setup satisfies Condition 3 because $r^{\star}\notin\mathcal{F}$: for any $\phi$, the constraint on $\psi$ implies $r_{\phi}(a_{3})=\tfrac{1}{2}(r_{\phi}(a_{1})+r_{\phi}(a_{2}))$ so $r_{\phi}(a_{3})=r$ whenever $r_{\phi}(a_{1})=r_{\phi}(a_{2})=r$, contradicting the ground-truth reward $r^{\star}(a_{3})=0$.
@@
where $\Delta:=\psi(a_{1})-\psi(a_{2})$. This expression is minimized at $\Delta^{\top}\phi=0$, which implies $r_{\phi}(a_{1})=r_{\phi}(a_{2})$ and $r_{\phi}(a_{3})=r_{\phi}(a_{1})$, i.e., the learned reward is constant: $r_{\textup{RLHF}}(a)=C$ for all $a\in\mathcal{Y}$.
@@ C.5 Proof of Proposition 6
C.6 Proof of Proposition 8
Construction 1: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. We first construct an environment satisfying Condition 6 such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. Consider the same setup as in Section C.4, but define the policy class as $\Pi=\Delta(\mathcal{Y})\setminus\{\pi^{\star}\}$, where $\pi^{\star}$ is the optimal policy under $r^{\star}$. This ensures that $\pi^{\star}\notin\Pi$, while $\mathcal{F}\subset\mathcal{F}_{\Pi}$, satisfying Condition 6.
As shown in Section C.4, RLHF learns a constant reward model and returns the uniform policy $\pi_{\textup{RLHF}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$. In contrast, DPO directly optimizes policy parameters from preference data and can converge to a policy arbitrarily close to $\pi^{\star}$, which lies on the boundary of $\Pi$. This yields a strict sub-optimality gap:
@@ C.6 Proof of Proposition 8
The corresponding surrogate reward class is $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\hat{r}_{\theta}(a)=\beta\theta^{\top}\psi(a),\;\theta\in\mathbb{R}^{2}\}$. We now define a strictly smaller reward model class =$\mathcal{F}=\{\hat{r}_{\theta_{R}}\}$ where
@@
In contrast, DPO learns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, as shown in Section C.3. Its value is:
@@ C.7 Proof of Proposition 9
From Section C.3, we know that DPO learns a constant reward model under this feature structure and returns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
@@
[…]
@@
C.8 Numerical Proof of Proposition 4
@@
C.9 Numerical Proof of Proposition 7
By Proposition 6, we have $\pi_{\textup{RLHF}}=\pi_{\textup{DPO}}$. Now we only need to construct an environment under Condition 5, such that online DPO can outperform offline DPO. We can borrow the whole setting in Section C.8, while resetting the ground-truth reward as:
@@ C.9 Numerical Proof of Proposition 7
C.10 Proofs of Theorems 10 and 11
Recall that to train a (surrogate) reward model, people first collect a dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then ask human annotators to label these pairs to get a human preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{N}$. Following BT model, $y_{1}$ is preferred over $y_{2}$, (i.e. $y_{w}=y_{1}$ and $y_{l}=y_{2}$), w.p. $\sigma(r^{\star}(y_{1})-r^{\star}(y_{2})$, where $r^{\star}(y)=(\theta^{\star})^{\top}\psi(y)$, $\theta^{\star}\in\mathbb{R}_{+}$ is the ground-truth reward vector, $\psi(y)$ is the feature vector satisfying $\|\psi(y)\|_{2}\leq L$, and $L\in\mathbb{R}_{+}$. The MLE estimator is defined as:
$\hat{\theta}_{\textup{MLE}}=\underset{\theta\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\theta^{\top}(\psi(y_{w})-\psi(y_{l})))\;,$ (6)
where $\Theta_{B}=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $B\in\mathbb{R}_{+}$. And we assume $\theta^{\star}\in\Theta_{B}$. The data-dependent performance measure is the data-induced semi-norm [Zhu et al., 2023], which is defined as:
@@ C.10 Proofs of Theorems 10 and 11
Remark 5.
Note that $\hat{\theta}_{1}$ has been fixed as $e_{1}$ in our context, we thus only need to measure $\hat{\theta}_{0}$.
Before proceeding, we first prepare some ingredients. It is worth noting that those lemmas below only work for the one-token scenario, and we will adopt them in the dual-token prediction task later.
@@
For a sample size $n\geq c_{1}\text{tr($\Sigma_{\mathcal{D}}$)}$, any estimator $\hat{\theta}$ based on $n$ samples has a lower bound as:
@@
Here $c_{1}$ is a numerical constant. This lemma is to establish a information-theoretical lower bound of one-token reward learning. Therefore, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ defined in Equation 6 is lower bounded.
Lemma 2 (Lemma 3.1 of Zhu et al. [2023]).
With probability at least $1-\delta$, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ has an upper bound:
@@
Definition 3 ($\ell_{0}$-constrained estimator).
$\hat{\theta}_{\ell_{0}}^{k}\in\underset{\theta\in\Theta_{B},\|\theta\|_{0}\leq k}{\operatorname{argmin}}\;\mathcal{L}_{\textup{MLE}}(\theta)\;.$
Lemma 3 (Theorem 3.2 of Yao et al. [2025]).
Consider $\|\theta^{\star}\|_{0}=k$ and $k\leq d/2$, then with probability at least $1-\delta$, the $\ell_{0}$-constrained estimator $\hat{\theta}_{\ell_{0}}^{k}$ has an upper bound:
$\|\hat{\theta}_{\ell_{0}}^{k}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{k\log(d/k)+\log(1/\delta)}{n}\right)\;.$
@@
Consider $\|\theta^{\star}\|_{0}=k$, then with probability at least $1-\delta$, the $\ell_{1}$-regularized estimator $\hat{\theta}_{\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\sqrt{\frac{\log 2d+\log(1/\delta)}{n}}\right)$ has an upper bound:
$\|\hat{\theta}_{\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\sigma)}{n}}\right)\;.$
@@
Consider $\tau\in\Theta_{B}$, $\|\theta^{\star}-\tau\|_{0}=k$, then with probability at least $1-\delta$, the relative $\ell_{1}$-regularized estimator $\hat{\theta}_{\textup{rel}\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\frac{\log 2d+\log(1/\delta)}{n}\right)$ has an upper bound:
@@
$\|\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
Proof of this lemma is deferred to Section C.11.
Proof of Theorem 10. We first prepare the settings and assumptions. For the feature mapping, we set $\psi(y,\omega)=\psi(\omega)$. For DTSP task, we first randomly collect a one-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{N}$, and ask human annotators to label these pairs. Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)}y_{w}^{(i)},y_{l}^{(i)}y_{l}^{(i)}\}_{i=1}^{N}$. And we further assume the Gram matrix $\frac{1}{n}\sum_{i=1}^{n}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}$ to be full-rank, and $(\textbf{r}_{\textup{sparse}})_{1}=-0.5$. Then by manipulating $\pi_{{\mathsf{ref}}}$, we can set:
$\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(\omega)_{1})=\left(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}}\right)^{\top}\psi(y)+C_{5}\;,$
where $\textbf{r}_{\textup{dense}}\in\mathbb{R}^{d}$ is a dense vector and $C_{5}\in\mathbb{R}$ is an offset. And this is realizable as long as:
$\max_{\omega}\psi(\omega)_{1}-\min_{\omega}\psi(\omega)_{1}\geq\max_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)-\min_{y}(\textbf{r}_{\textup{dense}}-\textbf{r}_{\textup{sparse}})^{\top}\psi(y)\;,$
which can be taken as a mild assumption on the features because $\psi$ and $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}}$ are all bounded. Now we are ready to prove the result.
Recall that:
@@
$(\theta_{p,0}^{\star})^{\top}\psi(y)=\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(r^{\star}(y,\omega)/\beta)$ $+C_{4}=\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\log\underset{\omega\sim\pi_{{\mathsf{ref}}}(\cdot|y)}{\mathbb{E}}\exp(\psi(\omega)_{1})+C_{4}\;,$
we thus have $\theta^{\star}_{r,0}=\textbf{r}_{\textup{sparse}}$ while $\theta^{\star}_{p,0}=\textbf{r}_{\textup{dense}}$, due to the non-singularity of the Gram matrix.
Then observe the empirical MLE objective of pure reward learning:
$\hat{\theta}_{r}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{r,0}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))\;,$
and the empirical MLE objective of surrogate reward learning:
$\hat{\theta}_{p}$ $=\underset{\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\hat{r}_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-\hat{r}_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{p,0}$ $=\underset{\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$
We can thus directly apply Lemma 2 to obtain that w.p. $1-\delta$, we have:
$\|\hat{\theta}_{p,0}-\theta^{\star}_{p,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;,$
and
$\|\hat{\theta}_{r,0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{d+\log(1/\delta)}{n}\right)\;.$
Note that $(\textbf{r}_{\textup{sparse}})_{1}=-0.5$, we can have an $\ell_{0}$-constrained estimator for the reward model:
$\hat{\theta}_{r,\ell_{0}}^{k}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\|\theta_{0}+e_{1}\|_{0}\leq k,\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))\;,$
$\implies\hat{\theta}_{r,\ell_{0},0}^{k}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\|\theta_{0}+e_{1}\|_{0}\leq k}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))\;,$
then Lemma 3 implies that w.p. $1-\delta$,
$\|\hat{\theta}_{r,\ell_{0},0}^{k}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\frac{k\log(d/k)+\log(1/\delta)}{n}\right)\;.$
And we also have an $\ell_{1}$-regularized estimator:
$\hat{\theta}_{r,\ell_{1}}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(r_{\theta}(y_{w}^{(i)}y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}y_{l}^{(i)}))+\gamma\|\theta_{0}+e_{1}\|_{1}\;,$
$\implies\hat{\theta}_{r,\ell_{1},0}$ $=\underset{\theta_{0}+e_{1}\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{N}\sum_{i=1}^{N}\log\sigma(\beta(\theta_{0}+e_{1})^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)}))+\gamma\|\theta_{0}+e_{1}\|_{1}\;.$
then Lemma 4 implies that w.p. $1-\delta$,
$\|\hat{\theta}_{r,\ell_{1},0}-\theta^{\star}_{r,0}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(2d)+k\log(1/\delta)}{n}}\right)\;.$
While for the surrogate reward model, note that $\hat{r}^{\star}_{p,0}$ is dense, and $e_{1}-\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}$ is intractable, it thus cannot apply $\ell_{0}$-constrained or $\ell_{1}$-regularized estimator to decrease the sample complexity.
@@
Proof of Theorem 11. We set $\psi(y,\omega)=\psi(\omega)+(\textbf{r}_{\textup{dense}}^{\top}\psi(y))e_{1}$, where $\textbf{r}_{\textup{dense}}\in\mathbb{R}^{d}$ is a dense vector. For DTSP task, we first randomly collect a one-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{N}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{N}$, and ask human annotators to label these pairs. Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)}y_{w}^{(i)},y_{l}^{(i)}y_{l}^{(i)}\}_{i=1}^{N}$. And we further assume the Gram matrix $\frac{1}{n}\sum_{i=1}^{n}(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))(\psi(y_{w}^{(i)})-\psi(y_{i}^{(i)}))^{\top}$ to be full-rank. Then by manipulating $\pi_{{\mathsf{ref}}}$, we can set:

### v2

@@ B.3 Proof of Proposition 3
$\mathcal{L}_{\textup{DPO}}(\pi_{\theta})$ $=-\sum_{a\neq a^{\prime}}\left[\sigma(r^{\star}(a)-r^{\star}(a^{\prime}))\log\sigma\left(\beta\,\theta^{\top}(\psi(a)-\psi(a^{\prime})\right)\right]$
@@ B.4 Proof of Proposition 5
For the strict gap, consider a multi-armed bandit setting analogous to Section B.3: first, define the action space $\mathcal{Y}=\{a_{1},a_{2},a_{3}\}$. Let the ground-truth reward function satisfy:
@@
The key difference from the earlier construction lies in the choice of model classes. We define: the linear reward class $\mathcal{F}=\{r_{\phi}:\phi\in\mathbb{R}^{d}\}$ by $r_{\phi}(a):=\phi^{\top}\psi(a)$, and the policy class $\Pi=\Delta(\mathcal{Y})$ with reference policy $\pi_{{\mathsf{ref}}}=\textrm{Unif}(\mathcal{Y})$. This setup satisfies Condition 3 because $r^{\star}\notin\mathcal{F}$: for any $\phi$, the constraint on $\psi$ implies $r_{\phi}(a_{3})=\tfrac{1}{2}(r_{\phi}(a_{1})+r_{\phi}(a_{2}))$ so $r_{\phi}(a_{3})=r$ whenever $r_{\phi}(a_{1})=r_{\phi}(a_{2})=r$, contradicting the ground-truth reward $r^{\star}(a_{3})=0$.
@@
where $\Delta:=\psi(a_{1})-\psi(a_{2})$. This expression is minimized maximized at $\Delta^{\top}\phi=0$, which implies $r_{\phi}(a_{1})=r_{\phi}(a_{2})$ and $r_{\phi}(a_{3})=r_{\phi}(a_{1})$, i.e., the learned reward is constant: $r_{\textup{RLHF}}(a)=C$ for all $a\in\mathcal{Y}$.
@@ B.5 Proof of Proposition 6
B.6 Proof of Proposition 8
Construction 1: $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. We first construct an environment satisfying Condition 6 such that $V_{r^{\star}}^{\pi_{\textup{RLHF}}}<V_{r^{\star}}^{\pi_{\textup{DPO}}}$. Consider the same setup as in Section B.4, but define the policy class as $\Pi=\Delta(\mathcal{Y})\setminus\{\pi^{\star}\}$, where $\pi^{\star}$ is the optimal policy under $r^{\star}$. This ensures that $\pi^{\star}\notin\Pi$, while $\mathcal{F}\subset\mathcal{F}_{\Pi}$, satisfying Condition 6.
As shown in Section B.4, RLHF learns a constant reward model and returns the uniform policy $\pi_{\textup{RLHF}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$. In contrast, DPO directly optimizes policy parameters from preference data and can converge to a policy arbitrarily close to $\pi^{\star}$, which lies on the boundary of $\Pi$. This yields a strict sub-optimality gap:
@@ B.6 Proof of Proposition 8
The corresponding surrogate reward class is $\mathcal{F}_{\Pi}=\{\hat{r}_{\theta}:\hat{r}_{\theta}(a)=\beta\theta^{\top}\psi(a),\;\theta\in\mathbb{R}^{2}\}$. We now define a strictly smaller reward model class $\mathcal{F}=\{\hat{r}_{\theta_{R}}\}$ where
@@
In contrast, DPO learns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, as shown in Section B.3. Its value is:
@@ B.7 Proof of Proposition 9
From Section B.3, we know that DPO learns a constant reward model under this feature structure and returns the uniform policy $\pi_{\textup{DPO}}=\pi_{{\mathsf{ref}}}$, independent of $r$ and $\beta$.
@@
$=\frac{e^{x/2}+e^{-x/2}}{Z(x)}-\beta\left[\frac{e^{x/2}}{Z(x)}\log\left(\frac{e^{x/2}}{Z(x)}\right)+\frac{e^{-x/2}}{Z(x)}\log\left(\frac{e^{-x/2}}{Z(x)}\right)+\frac{1}{Z(x)}\log\left(\frac{1}{Z(x)}\right)\right]+\text{(constant)}\;,$
@@
B.8 Numerical Proof of Proposition 4
Since the exact solution for online DPO is hard to compute, we didn’t find elegant proofs for these two propositions. They are examined correct numerically.
@@ B.8 Numerical Proof of Proposition 4
B.9 Numerical Proof of Proposition 7
By Proposition 6, we have $\pi_{\textup{RLHF}}=\pi_{\textup{DPO}}$. Now we only need to construct an environment under Condition 5, such that online DPO can outperform offline DPO. We can borrow the whole setting in Section B.8, while resetting the ground-truth reward as:
@@ B.9 Numerical Proof of Proposition 7
B.10 Formal Statement of Theorem 10 and Proof
B.10.1 Preliminaries of single-token prediction
Before proceeding, we first prepare some ingredients for the single-token prediction task.
Basic setting. Recall that to train a (surrogate) reward model, people first collect a dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then ask human annotators to label these pairs to get a human preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$. Following BT model, $y_{1}$ is preferred over $y_{2}$, (i.e. $y_{w}=y_{1}$ and $y_{l}=y_{2}$), w.p. $\sigma(r^{\star}(y_{1})-r^{\star}(y_{2}))$, where $r^{\star}(y)=(\theta^{\star})^{\top}\psi(y)$, $\theta^{\star}\in\mathbb{R}_{+}$ is the ground-truth reward vector, $\psi(y)$ is the feature vector satisfying $\|\psi(y)\|_{2}\leq L$, and $L\in\mathbb{R}_{+}$. The MLE estimator is defined as:
$\hat{\theta}_{\textup{MLE}}\in\underset{\theta\in\Theta_{B}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(\theta^{\top}(\psi(y_{w}^{(i)})-\psi(y_{l}^{(i)})))\;,$ (8)
where $\Theta_{B}=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $B\in\mathbb{R}_{+}$. And we assume $\theta^{\star}\in\Theta_{B}$. The empirical performance measure is the data-induced semi-norm (see, e.g., [Zhu et al., 2023]), defined as:
@@ B.10.1 Preliminaries of single-token prediction
And we assume $\Sigma_{\mathcal{D}}$ to be non-singular.
Note that the lemmas below only work for the single-token scenario, and we will adopt them in the dual-token prediction task later. The results quoted below from [Yao et al., 2025] follow directly from a long line of work on compressed sensing and sparse recovery based on restricted isometry (or restricted eigenvalue) properties [Candes et al., 2006], recast for the preference learning setting.
@@
For a sample size $n\geq c_{1}\text{tr($\Sigma_{\mathcal{D}}^{-1}$)}$, any estimator $\hat{\theta}$ based on $n$ samples has a lower bound as:
@@
Here $c_{1}$ is a constant independent of data. This lemma is to establish an information-theoretical lower bound for single-token reward learning.
Lemma 2 (Lemma 3.1 of Zhu et al. [2023]; see also Shah et al. [2015]).
W.p. at least $1-\delta$, the estimation error of the MLE estimator $\hat{\theta}_{\textup{MLE}}$ has an upper bound:
@@
Consider $\|\theta^{\star}\|_{0}=k$, then w.p. at least $1-\delta$, the $\ell_{1}$-regularized estimator $\hat{\theta}_{\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\sqrt{\frac{\log(d)+\log(1/\delta)}{n}}\right)$ has an upper bound:
$\|\hat{\theta}_{\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
@@
Consider $\tau\in\Theta_{B}$, $\|\theta^{\star}-\tau\|_{0}=k$, then w.p. at least $1-\delta$, the relative $\ell_{1}$-regularized estimator $\hat{\theta}_{\textup{rel}\ell_{1}}$ with an appropriate $\gamma=\Theta\left(\frac{\log(d)+\log(1/\delta)}{n}\right)$ has an upper bound:
@@
$\|\hat{\theta}_{\textup{rel}\ell_{1}}-\theta^{\star}\|_{\Sigma_{\mathcal{D}}}^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;.$
Proof of this lemma is given in Section B.10.4.
B.10.2 Formal statement of Theorem 10
Assumption 11 (Task configuration).
@@ B.10.2 Formal statement of Theorem 10
Recall that in DTSP task, we have $r^{\star}(y,\omega)=\beta\textbf{r}_{\textup{sparse}}^{\top}\psi(y)+\beta e_{1}^{\top}\psi(y,\omega)$, where $\psi(y),\psi(y,\omega),\textbf{r}_{\textup{sparse}}\in\mathbb{R}_{d}$, and $\|\textbf{r}_{\textup{sparse}}\|_{0}=k$, $k\ll d$. We further assume $B,L\in\mathbb{R}_{+}$, $\Theta_{B}:=\{\theta\in\mathbb{R}_{d}:\|\theta\|_{2}\leq B\}$, $\textbf{r}_{\textup{dense}},\textbf{r}_{\textup{sparse}},e_{1}+\textbf{r}_{\textup{dense}}+\textbf{r}_{\textup{sparse}}\in\Theta_{B}$, $\|\psi(y)\|_{2}\leq L$, and $\psi(y,\omega)=\psi(\omega)+(\textbf{r}_{\textup{dense}}^{\top}\psi(y))e_{1}$.
Assumption 12 (Preference data collection).
@@
For DTSP task, we first collect a single-token dataset $\mathcal{D}^{\dagger}=\{y_{1}^{(i)},y_{2}^{(i)}\}_{i=1}^{n}$, and then duplicate it as $\mathcal{D}^{\ddagger}=\{y_{1}^{(i)}y_{1}^{(i)},y_{2}^{(i)}y_{2}^{(i)}\}_{i=1}^{n}$, and ask human annotators to label these pairs. Now we have collected a dual-token preference dataset $\mathcal{D}=\{y_{w}^{(i)},y_{l}^{(i)}\}_{i=1}^{n}$, where $y_{w}^{(i)}=\bar{y}_{w}^{(i)}\bar{y}_{w}^{(i)}$ and $y_{l}^{(i)}=\bar{y}_{l}^{(i)}\bar{y}_{l}^{(i)}$. And we further assume that the Gram matrix $\Sigma_{\mathcal{D}}:=\frac{1}{n}\sum_{i=1}^{n}(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))(\psi(\bar{y}_{w}^{(i)})-\psi(\bar{y}_{l}^{(i)}))^{\top}$ is non-singular, and $\text{tr}(\Sigma_{\mathcal{D}}^{-1})=\mathcal{O}(d)$.
Theorem 13 (Formal separation theorem).
Under token-level linear parameterization and Assumptions 11 and 12, there exists an environment for DTSP task, s.t. by estimating from a preference dataset $\mathcal{D}$ with $n$ samples under $\theta_{1}=e_{1}$ constraint, the estimation error of the reward model $\hat{\theta}_{r}$ can be reduced to $\tilde{\mathcal{O}}(\sqrt{k\log d/n})$ using a (computationally efficient) $\ell_{1}$-regularized estimator:
$\hat{\theta}_{r,\textup{rel}\ell_{1}}$ $\in\underset{\theta_{0}+e_{1}+\textbf{r}_{\textup{dense}}\in\Theta_{B},\theta_{1}=e_{1}}{\operatorname{argmin}}-\frac{1}{n}\sum_{i=1}^{n}\log\sigma(r_{\theta}(y_{w}^{(i)})-r_{\theta}(y_{l}^{(i)}))+\gamma\|\theta_{0}\|_{1}\;,$
i.e., w.p. $1-\delta$,
$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{w}^{(i)})-r_{\hat{\theta}_{r,\textup{rel}\ell_{1}}}(y_{l}^{(i)}))\right]^{2}=\mathcal{O}\left(\sqrt{\frac{k\log(d)+k\log(1/\delta)}{n}}\right)\;,\;$
while the estimation error of any estimator for the DPO model $\hat{\theta}_{p}$ is lower bounded by $\Omega(d/n)$:
$\frac{1}{n}\sum_{i=1}^{n}\left[(r^{\star}(y_{w}^{(i)})-r^{\star}(y_{l}^{(i)}))-(r_{\hat{\theta}_{p}}(y_{w}^{(i)})-r_{\hat{\theta}_{p}}(y_{l}^{(i)}))\right]^{2}=\Omega\left(\frac{d}{n}\right)\;.$
B.10.3 Proof of Theorem 13
Let $\pi_{{\mathsf{ref}}}(\cdot|y)$ be identical for all $y$, then we have